Find the ratio of strain energies stored in bars A and B of the same axial tensile loads. The bar A is of 50mm diameter through out its length, while bar B though of same length as of A but has diameter of 25 mm for the middle one - third of its length and the remainder is of 50 mm . Assume gradual load
1
Expert's answer
2020-12-15T11:46:34-0500

The strain energy formula


"U=F\delta/2",


where "\delta=" compression, "F=" force applied. So,


"U_2/U_1=(F\delta_2/2)/(F\delta_1/2)=\delta_2/\delta_1."


"\delta_1=\frac{FL}{A_1E}" and "\delta_2=\frac{F(1/3)L}{A_2E}+\frac{F(1/3)L}{A_1E}+\frac{F(1/3)L}{A_1E}="


"=\frac{F(1/3)L}{A_2E}+\frac{F(2/3)L}{A_1E}."


"U_2/U_1=\delta_2/\delta_1=\frac{\frac{F(1/3)L}{A_2E}+\frac{F(2/3)L}{A_1E}}{\frac{FL}{A_1E}}=\frac{A_1+2A_2}{3A_2}="


"=\frac{\pi d_1^2/4+2\pi d_2^2/4}{3\pi d_2^2/4}=\frac{d_1^2+2d_2^2}{3d_2^2}=\frac{0.05^2+2\cdot0.025^2}{3\cdot0.025^2}=2(J)"